OpenStudy (anonymous):

tan^2 x + 4 = 2sec^2 x + tanx

OpenStudy (anonymous):

You have to prove the identity? @Yellowpanda

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

Solve for x?

OpenStudy (anonymous):

yes xD

OpenStudy (anonymous):

Well the obvious approach here would be to benefit here from the following identity:\[\sec^2(x)=\tan^2(x)+1\]

OpenStudy (anonymous):

Can you see yourself using that identity to solve for x? @Yellowpanda

OpenStudy (anonymous):

not really (

OpenStudy (anonymous):

Use that identity to replace sec^2(x), then bring the equation on one side. That way you'll have a quadratic so solve for and that will lead you to the answers for x.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

okay lemme try that

OpenStudy (anonymous):

\[\sec^2(x)=\tan^2(x) +1 \rightarrow \tan^2(x)+4=2(\tan^2(x)+1)+\tan(x)\]\[\rightarrow \tan^2(x)+4=2\tan^2(x)+2+\tan(x) \rightarrow \tan^2(x) + \tan(x) -2=0\]Now that we have a quadratic, can you solve for x? @Yellowpanda

OpenStudy (anonymous):

\[Tan ^{2}x+4=2\sec ^{2}x + Tanx\] Applying the formula Genius gave, \[Tan ^{2}x + 4 =2(\tan ^{2}x+1) + Tanx\]\[Tan ^{2}x + 4 = 2Tan ^{2}x + 2 + Tanx\] \[0=2Tan ^{2}x-Tan ^{2}x +2 - 4 + Tanx\] \[0 = Tan ^{2}x -2 + Tanx\] \[0 = Tan ^{2}x + Tanx - 2\]

OpenStudy (anonymous):

yes XD i got it

OpenStudy (anonymous):

thank you guys!!

OpenStudy (anonymous):

Good job! :D @Yellowpanda

OpenStudy (anonymous):

@Yellowpanda Give @genius12 a medal.